OpenStudy (anonymous):

What are the foci of the hyperbola given by the equation 16x2 – 9y2 + 64x – 72y – 224 = 0 ?

OpenStudy (anonymous):

i put (–3, 4) and (7, 4)

OpenStudy (tkhunny):

What did you get when you completed the square in x and in y?

OpenStudy (anonymous):

16x²+64x+1024-9y²-72y+1296=224

OpenStudy (tkhunny):

That was a little rearranging. How about the Completing the Square part? It should end up like this: \(16(x-Something)^{2} - 9(y-something)^{2} = Constant^{2} \)

OpenStudy (tkhunny):

BTW - You added 1024 and 1296 to one side of the equation. Does this make sense to you?

OpenStudy (anonymous):

oh yes i forgot that part. 16x²+64x+1024-9y²-72y+1296=224+1024+1296 16(x²+4x+64)-9(y²-8y+144)=2544

OpenStudy (anonymous):

wait did i do something wrong ?

OpenStudy (tkhunny):

\(16x^{2} + 64x = 16(x^{2} + 4x) = 16(x^{2} + 4x + 4 - 4) = 16(x^{2} + 4x + 4) - 16(4)\) = \(16(x^{2} + 4x + 4) - 64 = 16(x+2)^{2} - 64\) Not sure how you managed "64" inside the parentheses.

OpenStudy (anonymous):

okk so whats the next step to get the answer ?

OpenStudy (tkhunny):

Do the same thing for 'y'. You seem to have the right idea. Just be a little more careful.

OpenStudy (anonymous):

9(y-2)²-144 ?

OpenStudy (tkhunny):

Not quite. \(-9y^{2} - 72y =\) \(-9(y^{2} + 8y) =\) \(-9(y^{2} + 8y + 16 - 16) =\) \(-9(y^{2} + 8y + 16) - (-9)(16)\) \(-9(y+4)^{2} + 144\) You are SO CLOSE!! Keep working on it. You'll get it.

OpenStudy (anonymous):

ooppss :( . Thankyouu.(: And the next step is ?

OpenStudy (tkhunny):

Almost done. Just put it all together.

OpenStudy (anonymous):

16(x+2)² - 64 9(y+4)²+144

OpenStudy (anonymous):

16(x+2)² - 64 + 9(y+4)²+144 like this ?

OpenStudy (tkhunny):

\( = 224 + 64 - 144\)

OpenStudy (anonymous):

144

OpenStudy (tkhunny):

Sorry, I was a little confused. Result should be \(16(x+2)^{2} + 9(y+4)^{2} = 144\) Now what?

OpenStudy (anonymous):

umm 16+9 ?

OpenStudy (tkhunny):

What is that? Divide by 144 and see if anything nice happens.

OpenStudy (anonymous):

16+9=25 but it doesnt go into 144 evenly

OpenStudy (tkhunny):

I'm still wondering why you did that. \(\dfrac{(x+2)^2}{9} - \dfrac{(y+4)^{2}}{16} = 1\) or \(\dfrac{(x+2)^2}{3^{2}} - \dfrac{(y+4)^{2}}{4^{2}} = 1\)

OpenStudy (anonymous):

wait so , ohh nvm okk. so now what do i do ?

OpenStudy (tkhunny):

Once you find this wonderful form, you just start picking off information. Center: ?? a = ?? Location of vertices: ?? b = ?? Location of fake vertices: ?? Equations of asymptotes: ?? c = ?? Location of Foci: ?? Start filling in the blanks.

OpenStudy (anonymous):

0.o. Ughh idk what im doing. i give up :(

OpenStudy (tkhunny):

Just one thing at a time. Center: (-2,-4) -- Do you see where this comes from?

OpenStudy (anonymous):

yhes from the parenthesis . just the opposite.

OpenStudy (tkhunny):

The standard form is x-h and y-k in there. So x + 2 = x - (-2) and there is the -2. Now 'a' is half the length of the major axis. a = 3 -- Do you see where that comes from?

OpenStudy (anonymous):

yes i get it. so b i y-(-4) wich is 5 ?

OpenStudy (anonymous):

or maybe not .

OpenStudy (tkhunny):

How did 4 magically turn into 5? The center is (-2,-4) a = 3 (half the length of the major axis) b = 4 (half the length of the phantom axis) Those are from the denominators.

OpenStudy (anonymous):

becausee you made x - (-2) =3 so i thought y-(-4) = 5 . oohh ok nvm.

OpenStudy (anonymous):

hello ?

OpenStudy (tkhunny):

I do not understand what you are doing. x + 2 leads to h = -2 y + 4 leads to k = -4 This means the center is at (-2,-4). This has nothing to do with a or b or the c, yet to come. Do we have the center, yet?